# Deviations from the Nash equilibrium and emergence of tacit collusion in a two-player optimal execution game with reinforcement learning

### v1

@@ Abstract
The use of reinforcement learning algorithms in financial trading is becoming increasingly prevalent. However, the autonomous nature of these algorithms can lead to unexpected outcomes that deviate from traditional game-theoretical predictions and may even destabilize markets. In this study, we examine a scenario in which two autonomous agents, modeled with Double Deep Q-Learning, learn to liquidate the same asset optimally in the presence of market impact, using the Almgren-Chriss (2000) framework. Our results show that the strategies learned by the agents deviate significantly from the Nash equilibrium of the corresponding market impact game. Notably, the learned strategies exhibit tacit collusion, closely aligning with the Pareto-optimal solution. We further explore how different levels of market volatility influence the agents’ performance and the equilibria they discover, including scenarios where volatility differs between the training and testing phases.
@@ 1 Introduction
The increasing automation of trading over the past three decades has profoundly transformed financial markets. The availability of large, detailed datasets has facilitated the rise of algorithmic trading—a sophisticated approach to executing orders that leverages the speed and precision of computers over human traders. By drawing on diverse data sources, these automated systems have revolutionized how trades are conducted. In 2019, it was estimated that approximately 92% of trading in the foreign exchange (FX) market was driven by algorithms as reported in [kissell2020algorithmic]. The rapid advancements in Machine Learning (ML) and Artificial Intelligence (AI) have significantly accelerated this trend, leading to the widespread adoption of autonomous trading algorithms. These systems, particularly those based on Reinforcement Learning (RL), differ from traditional supervised learning models. Instead of being trained on labeled input/output pairs, RL algorithms explore a vast space of potential strategies, learning from past experiences to identify those that maximize long-term rewards. This approach allows for the continuous refinement of trading strategies, further enhancing the efficiency and effectiveness of automated trading.
@@
The flexibility of RL comes with significant potential costs, particularly due to the opaque, black-box nature of these algorithms. This opacity can lead to unexpected outcomes that may destabilize the system they control—in this case, financial markets—through the actions they take, such as executing trades. The complexity and risk increase further when multiple autonomous agents operate simultaneously within the same market. The lack of transparency in RL algorithms may result in these agents inadvertently learning joint strategies that deviate from the theoretical Nash equilibrium, potentially leading to unintended market manipulation. Among the various risks, the emergence of collusion is particularly noteworthy. Even without explicit instructions to do so, RL agents may develop cooperative strategies that manipulate the market, a phenomenon known as tacit collusion. This type of emergent behavior is especially concerning because it can arise naturally from the interaction of multiple agents, posing a significant challenge to market stability and fairness.
@@
In this paper, we investigate the equilibria and the emergence of tacit collusion in a market where two autonomous agents, driven by deep RL, engage in an optimal execution task. While tacit collusion in RL settings has been explored in areas such as market making (as reviewed below), much less is known about how autonomous agents behave when trained to optimally trade a large position. We adopt the well-established Almgren-Chriss framework first introduced in [almgren], for modeling market impact, focusing on two agents tasked with liquidating large positions of the same asset. The agents are modeled using Double Deep Q-Learning (DDQL) and engage in repeated trading to learn the optimal equilibrium of the associated open-loop game.
@@
For this specific setup, the Nash equilibrium of the game has been explicitly derived in [schied2017state], providing a natural benchmark against which we compare our numerically derived equilibria. In addition, we explicitly derive the Pareto-optimal strategy and numerically characterize the Pareto-efficient set of solutions. Our primary goal is to determine whether the two agents, without being explicitly trained to cooperate or compete, can naturally converge to a collusive equilibrium, given the existence and uniqueness of the Nash equilibrium.
@@
Our findings reveal that the strategies learned by the RL agents deviate significantly from the Nash equilibrium of the market impact game. Across various levels of market volatility, we observe that the learned strategies exhibit tacit collusion, closely aligning with the Pareto-optimal strategy. Given that financial market volatility is time-varying, we also examine the robustness of these strategies when trained and tested under different volatility regimes. Remarkably, we find that the strategies learned in one volatility regime remain collusive even when applied to different volatility conditions, underscoring the robustness of the training process.
@@
Optimal execution has been extensively studied in the financial literature. Starting from the seminal contributions in [bertsimas] and in [almgren], many authors have contributed to extend the model’s consistency with reality. Notable examples in this area are: [bouchaud2003fluctuations, bouchaud2009markets, gueant2012optimal, gatheral2012transient, obizhaeva2013optimal, gueant2015general, cartea2015algorithmic, cartea2016incorporating, casgrain2019trading]. In its basic setting, the optimal execution problem considers just one agent unwinding or acquiring a quantity $q_{0}$ of assets within a certain time window. However, if many other agents are considered to be either selling or buying, thus pursuing their own optimal execution schedules in the same market, the model becomes more complicated since it now requires modelling other agents’ behaviour the same market. Using a large dataset of real optimal execution, [bucci2020co] shows that the cost of an execution strongly depends on the presence of other agents liquidating the same asset. The increased complexity when treating the problem in this way allows for more consistency with reality and, at the same time, opens the path for further studies on how the agents interact with each other in such a context.
@@
The optimal execution problem with $n$-agents operating in the same market has thus been studied under a game theoretic lens in both its closed-loop (in [carmona2011predatory, micheli2023closed]) and open-loop (in [brunnermeier2005predatory, carlin2007episodic, schoneborn2009liquidation, schied2017state, schied2019market]) formulations. In more recent works, rather than just optimal execution, liquidity competition is also analysed (as in [drapeau2019fbsde, neuman2023trading]) along with market making problems (as in [cont2021interbank, cont2022dynamics]).
@@
In recent times, with the advancements of machine learning techniques, the original optimal execution problem has been extensively studied using RL techniques. Among the many [^3] For a more comprehensive overview of the state of the art on RL methods in finance, please refer to [sun2023reinforcement] and [hambly2023recent], some examples in the literature on RL techniques applied to the optimal execution problem are found in [ning2021double, schnaubelt2022deep, macrì2024reinforcement].
@@
Applications of multi-agent RL to financial problems are not incredibly numerous if compared to those that study the problem in a single agent scenario. Optimal execution in a multi-agent environment is tackled in [karpe2020multi], where the authors analyse the optimal execution problem form the standpoint of a many agents environment and an RL agent trading on Limit Order Book. The authors test and develop a multi-agent Deep Q-learning algorithm able to trade on both simulated and real order books data. Their experiments converge to agents who adopt only the so-called Time Weighted Average Price (TWAP) strategies in both cases. On the other hand, still on optimal execution problem’s applications of multi-agent RL, the authors in [bao2019multi] analyse the interactions of many agents on their respective optimal strategies under the standpoint of cooperative competitive behaviours, adjusting the reward functions in the Deep Deterministic Policy Gradient algorithm used, in order to allow for either of the two phenomena, but still using the basic model introduced in [almgren], and modelling the interactions of the agents via full reciprocal disclosure of their rewards and not considering the sum of the strategies to be a relevant feature for the permanent impact in the stock price dynamics.
@@
The existence of collusion and the emergence of collusive behaviours are probably the most interesting behaviours based on agents interaction in a market, as these are phenomena that might naturally arise in markets even if no instruction on how to and whether to collude or not, are given to the modelled agents. The emerging of tacit collusive behaviours has been analysed in various contexts. One of the first interesting examples is found in [waltman2008q], where using Q-learning the authors show how competing producers in a Cournot oligopoly learn how to increase prices above Nash equilibrium by reducing production, thus firms learn to collude even if usually do not learn to make the highest possible joint profit. Hence the firms converge to a ‘collusive’ behaviour rather than to a ‘proper’ collusion equilibrium. Still in Cournot oligopoly, authors in [abada2023artificial] apply multi-agent Q-learning in electricity markets, the evidences of a collusive behaviour do still arise and the authors postulate that the collusive behaviour may arise from imperfect exploration of the agents in such a framework. Similarly, authors in [hettich2021algorithmic] conclude that the use of deep RL leads to collusive behaviours faster than simulated tabular Q-learning algorithms. For what concerns financial markets, the problem becomes more involved since the actors in market do not directly set the price in a ‘one sided’ way as is the case for production economies that are studied in the previously cited works. Among the many contributions that study the emergence of collusive behaviours in financial markets, in [xiong2022interactions] it is shown how tacit collusion arises between market makers, modelled using deep RL, in a competitive market. In [cartea2022algorithmic] the authors show how market making algorithms tacitly collude to extract rents, and this behaviour strictly depends on tick size and coarseness of price grid. Finally, in [cont2022dynamics] the authors use a multi-agent deep RL algorithm to model market makers competing in the same market, the authors show how competing market makers learn how to adjust their quotes, giving rise to tacit collusion by setting spread levels strictly above the competitive equilibrium level.
@@
The paper is organised as follows: Section 2 introduces the market impact game theoretical setting, Section 3 introduces the DDQL algorithm for the multi-agent optimal execution problem, Section 4 discusses the results for different parameter settings, and finally, Section 5 provides conclusions and outlines further research directions.
@@ 2 Market impact game setting
The setup of our framework is based on the seminal Almgren-Chriss model first introduced in [almgren] for optimal execution. In this setting, a single agent wants to liquidate an initial inventory of $q_{0}$ shares within a time window $[0,T]$, which, in the discrete-time setting, is divided into $N$ equal time increments of length $\tau=T/N$. The main assumption of the model is that the mid-price (or the efficient price) evolves according to a random walk with a drift depending on the traded quantity. Moreover, the price obtained by the agent differs from the mid-price by a quantity which depends on the quantity traded in the interval. More formally, let $S_{t}$ and $\tilde{S}_{t}$ be the mid-price and the price received by the agent at time $t$, and let $v_{t}$ be the number of shares traded by the agent in the same interval; then the dynamics is given by
@@
$\begin{split}{S}_{t}&=S_{t-1}-g\left({v_{t}}/{\tau}\right)\tau+\sigma\tau^{\frac{1}{2}}\xi\\ \tilde{S_{t}}&={S}_{t-1}-h\left({v_{t}}/{\tau}\right)\end{split}$ (1)
@@
$S_{t}$ evolves because of a diffusion part $\xi\sim\mathcal{N}(0,1)$ multiplied by the price volatility $\sigma$ and a drift part, termed permanent impact, $g_{t}\left({v_{t}}/{\tau}\right)$, assumed to be linear and constant: $g\left({v_{t}}/{\tau}\right)=\kappa{v_{t}}/{\tau}$. The price $\tilde{S}_{t}$ received by the agent is equal to the mid-price $S_{t}$ but impacted by a temporary impact term also assumed to be linear and constant in time: $h(v_{t}/{\tau})=\alpha{v_{t}}/{\tau}$.
@@
The aim of the agent is to unwind their initial portfolio maximizing the cash generated by their trading over the $N$ time steps. This objective can be rewritten in terms of Implementation Shortfall (IS) that is defined in the single agent case as:
$IS(\vec{v})=S_{0}q_{0}-\sum^{N}_{t=1}\tilde{S}_{t}v_{t}$ (2)
where $\vec{v}=(v_{1},...,v_{N})^{\prime}\in{\mathbb{R}}^{N}$ is the vector containing the traded quantity in each time step. The optimisation problem of the agent can be written as
$\min_{\vec{v}}\,\,\mathbb{E}[IS(\vec{v})]-{\lambda}\mathbb{V}[IS(\vec{v})]\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ s.t.\leavevmode\nobreak\ \leavevmode\nobreak\ \sum_{t=1}^{N}v_{t}=q_{0}$ (3)
@@
where $\lambda$ is the risk aversion parameter of the agent. Under the linearity assumption of the two impacts, the problem can be easily solved analytically. In the following, we are going to consider risk neutral agents ($\lambda=0$) and in this case the optimal trading schedule is the TWAP, i.e. $v_{t}=q_{0}/N$, where the trading velocity is constant.
@@
Now we consider two agents selling an initial inventory $q_{0}$ shares within the same time window time window $[0,T]$. The traded quantities by the two agents in the time step $t$ are indicated with $v^{(1)}_{t}$ and $v^{(2)}_{t}$ and $V_{t}=v^{(1)}_{t}+v^{(2)}_{t}$ is the total quantity traded in the same interval. The equations for the dynamics are
@@
$\begin{split}{S}_{t}&=S_{t-1}-g\left({V_{t}}/{\tau}\right)\tau+\sigma\tau^{\frac{1}{2}}\xi\\ \leavevmode\nobreak\ \leavevmode\nobreak\ \tilde{S_{t}}^{(k)}&={S}_{t-1}-h\left({v^{(k)}_{t}}/{\tau}\right)\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ k=1,2\end{split}$ (4)
i.e. the mid-price is affected by the total traded volume $V_{t}$, while the price received by each agents depend on the quantity she trades.
@@
Since more than one agent is optimising her trading and the cash received depends on the trading activity of the other agent, the natural setting to solve this problem is the one of game theory. Since each agent is not directly aware of the selling activity of the other agent and the two agents interact through their impact on the price process, the resulting problem is an open-loop game. The existence and uniqueness of a Nash equilibrium in such a symmetric open-loop game has been studied in [schied2017state] in the case of n different players and linear impact functions. First of all, the defined the Nash equilibrium as:
Definition 1 (Nash equilibrium ([schied2017state])).
@@
In an $n$ players game, where $n\in\mathbb{N}$, with $q^{(1)}_{0},\dots,q^{(n)}_{0}\in\mathbb{R}$ initial inventory holdings and $\lambda_{1},\dots,\lambda_{n}$ non-negative coefficients of risk aversion. A Nash equilibrium for mean-variance optimisation as in Eq. (3) is a collection of $\mathbf{q^{*}}=\{\vec{q}_{1}^{*},\dots,\vec{q}_{n}^{*}\}$ inventory holdings such that for each $k\in n,\leavevmode\nobreak\ {\vec{q}_{k}}\in\mathcal{A}_{\text{det}}$ the mean variance functional is minimised:
$\mathbb{E}[IS(\vec{q}_{k}|\mathbf{q^{*}_{-k}})]-{\lambda_{k}}\mathbb{V}[IS(\vec{q}_{k}|\mathbf{{q}^{*}_{-k}})]$
@@
for each agent $k$ [^4] Where $\mathcal{A}_{\text{det}}$ is the set of all admissible deterministic trading strategies, for a more formal definition see [schied2017state]. Where $\mathbf{{q}^{*}_{-k}}$ are the strategies of the other players minus the strategy of the $k$-agent that is being considered.
@@
Then they proved that there exists a unique Nash equilibrium for the mean-variance optimisation problem. The unique Nash equilibrium strategy $q^{*(k)}_{t}$, the Nash remaining inventory of agent $k$ at time $t$, for $n$ players is given as a solution to the second-order system of differential equations:

### v2

# Deviations from the Nash equilibrium in a two-player optimal execution game with reinforcement learning

@@ Abstract
The use of reinforcement learning algorithms in financial trading is becoming increasingly prevalent. However, the autonomous nature of these algorithms can lead to unexpected outcomes that deviate from traditional game-theoretical predictions and may even destabilize markets. In this study, we examine a scenario in which two autonomous agents, modelled with Double Deep Q-Learning, learn to liquidate the same asset optimally in the presence of market impact, under the Almgren-Chriss (2000) framework. We show that the strategies learned by the agents deviate significantly from the Nash equilibrium of the corresponding market impact game. Notably, the learned strategies exhibit supra-competitive solution, which might be compatible with a tacit collusive behaviour, closely aligning with the Pareto-optimal solution. We further explore how different levels of market volatility influence the agents’ performance and the equilibria they discover, including scenarios where volatility differs between the training and testing phases.
@@ 1 Introduction
The increasing automation of trading over the past three decades has profoundly transformed financial markets. The availability of large, detailed datasets has facilitated the rise of algorithmic trading—a sophisticated approach to executing orders that leverages the speed and precision of computers over human traders. Using diverse data sources, these automated systems have revolutionized the way trades are conducted. In 2019, it was estimated that approximately 92% of the trading in the foreign exchange (FX) market was driven by algorithms [undefad]. The rapid advancements in Machine Learning (ML) and Artificial Intelligence (AI) have significantly accelerated this trend, leading to an increasing adoption of autonomous trading algorithms and, as argued in [undefc], this poses problems from a legislative point of view when it comes to deciding whether and how regulators can hold accountable firms for malpractices due to algorithms’ misuse. In the landscape of autonomous trading systems, increasing interest, by both practitioners and academics alike, is being shown on those based on Reinforcement Learning (RL), especially deep RL. Deep RL algorithms differ from traditional supervised learning models. Instead of being trained on labelled input/output pairs, RL algorithms explore a vast space of potential strategies, learning from past experiences to identify those that maximise long-term rewards. This approach allows for the continuous refinement of trading strategies, further enhancing the efficiency and effectiveness of automated trading.
@@
The flexibility of deep RL comes with significant potential costs, particularly due to the opaque, black-box nature of these algorithms. This opacity can lead to unexpected outcomes that may destabilize the system they control—in this case, financial markets—through the actions they take, such as executing trades. The complexity and risk increase further when multiple autonomous agents operate simultaneously within the same market. The lack of transparency in RL algorithms may result in these agents inadvertently learning joint strategies that deviate from the theoretical Nash equilibrium, potentially leading to unintended market manipulation. Among the various risks, the emergence of collusion is particularly notable. Even without explicit instructions to do so, RL agents may develop cooperative strategies that manipulate the market, a phenomenon known as tacit collusion. This type of emergent behaviour is especially concerning because it can arise naturally from the interaction of multiple agents, posing a significant challenge to market stability and fairness.
@@
In this paper, we investigate the equilibria and the emergence of supra-competitive equilibria and their implications in terms of tacit collusive behaviour in a market where two autonomous agents, modelled by deep RL, engage in an optimal execution task. Although tacit collusion in RL settings has been explored in areas such as market making (as reviewed below), much less is known about how autonomous agents behave when trained to optimally trade a large position. We adopt the well-established Almgren-Chriss framework first introduced in [undefb], for modelling market impact, focusing on two agents tasked with liquidating large positions of the same asset. The agents are modelled using Double Deep Q-Learning (DDQL) and engage in repeated trading to learn the optimal equilibrium of the associated open-loop game.
@@
For this specific setup, the Nash equilibrium of the game has been explicitly derived in [undefam], providing a natural benchmark against which we compare our numerically derived equilibria. In addition, we explicitly derive the Pareto-optimal strategy and numerically characterize the Pareto-efficient set of solutions. Our primary goal is to determine whether the two agents, without being explicitly trained to cooperate or compete, can naturally converge to a non Nash equilibrium, when it existence and uniqueness is known.
@@
Our findings reveal that the strategies learned by the RL agents deviate significantly from the Nash equilibrium of the market impact game. Across various levels of market volatility, we observe that the learned strategies exhibit seemingly tacit collusive behaviour, closely aligning with the Pareto-optimal strategy. Given that financial market volatility is time-varying, we also examine the robustness of these strategies when trained and tested under different volatility regimes. Remarkably, we find that the strategies learned in one volatility regime remain supra-competitive, even when adopted in various volatility conditions, underscoring the robustness of the training process.
@@
Optimal execution has been extensively studied in the financial literature. Starting from the seminal contributions in [undefe] and in [undefb], many authors have contributed to extend the model’s consistency with reality. Notable examples in this area are [undefp, undefo, undefq] for optimal control perspective, [undefh, undefg, undefz, undefw, undefaj] for the transient impact extensions and [undefy] for a limit order generalisation of the Almgren and Chriss framework. In its basic setting, the optimal execution problem considers just one agent unwinding or acquiring a quantity $q_{0}$ of assets within a certain time window. However, if many other agents are considered to be either selling or buying, thus pursuing their own optimal execution schedules in the same market, the model becomes more complicated since it now requires modelling other agents’ behaviour the same market. Using a large dataset of real optimal executions, [undefj] shows that the cost of an execution strongly depends on the presence of other agents liquidating the same asset. The increased complexity when treating the problem in this way allows for more consistency with reality and, at the same time, opens the path for further studies on how the agents interact with each other in such a context.
@@
The optimal execution problem with $n$-agents operating in the same market has thus been studied under a game theoretic lens in both its closed-loop (in [undefm, undefag]) and open-loop (in [undefi, undefl, undefao, undefam, undefal]) formulations. In more recent works, rather than just optimal execution, liquidity competition is also analysed (as in [undefv, undefah]) along with market making problems (as in [undefr, undefs]).
@@
In recent times, with the advancements of machine learning techniques, the original optimal execution problem has been extensively studied using RL techniques. Among the many [^3] For a more comprehensive overview of the state of the art on RL methods in finance, please refer to [undefap] and [undefaa], some examples in the literature on RL techniques applied to the optimal execution problem are found in [undefai, undefan, undefaf].
@@
Applications of multi-agent RL to financial problems are scarce if compared to those that study the problem in a single agent scenario. Optimal execution in a multi-agent environment is tackled in [undefac], where the authors analyse the optimal execution problem form the standpoint of a many agents environment and an RL agent trading on Limit Order Book. The authors test and develop a multi-agent Deep Q-learning algorithm able to trade on both simulated and real order books data. Their experiments converge to agents who adopt only the so-called Time Weighted Average Price (TWAP) strategies in both cases. On the other hand, still on optimal execution problem’s applications of multi-agent RL, the authors in [undefd] analyse the interactions of many agents on their respective optimal strategies under the standpoint of cooperative competitive behaviours, adjusting the reward functions in the Deep Deterministic Policy Gradient algorithm used, in order to allow for either of the two phenomena, but still using the basic model introduced in [undefb], and modelling the interactions of the agents via full reciprocal disclosure of their rewards and not considering the sum of the strategies to be a relevant feature for the permanent impact in the stock price dynamics.
The existence of collusion and the emergence of collusive behaviours are probably the most interesting phenomena based on agents interaction in a market, as these are circumstances that might naturally arise in markets even if no instruction on how to and whether to collude or not, are given to the modelled agents.
One of the first contributions to collusion theory of autonomous algorithms has been put forth in [undefaq]. In this contribution the authors show empirically by means of Q-learning, how two firms in a production game with linear demand function, can learn collusive behaviour even if no direct communication and explicit punishment strategies are present. Hence, in this contribution the authors show how competing producers in a Cournot oligopoly learn to increase prices above Nash equilibrium by reducing production, so that firms learn to collude even if they do not learn to obtain the highest possible joint profit. Thus, the firms reach a supra-competitive equilibrium, comparable to the collusive one even if no punishment mechanism is explicitly present or direct communication between the players is present. This remarkable result has been further analysed and extended in [undefk]. Here, the authors show how using Q-learning, two agents in a duopoly are able to retrieve supra-competitive equilibria without information exchange. The equilibrium is sustained by collusive strategies enforced by eventual punishment phases in the case where an agent deviates from said collusive equilibrium.
Notwithstanding the empirical emergence of collusive behaviour, some concerns have emerged on whether or not the supra-competitive outcomes of such games are indeed collusions or are due to some algorithm-specific features and thus, if these possibly tacit collusive behaviours have to be a major concern for the regulators. This topic has been investigated in [undef], where the authors consider battery and electricity markets to empirically prove how such agents would indeed punish any deviation from a collusive strategy by inducing shocks to the environment where the agents are playing after they have converged in the training phase. They conjecture that this seeming collusion is due to imperfect exploration of states and parameter initialisation. More recently, [undefa] expands this point of view by showing how in a Q-learning environment seemingly supra-competitive equilibria are due to the poor exploration policy of unsophisticated algorithms. They show how for simple Q-learning algorithms there is an auto-induced cooperative trap if states are not thoroughly explored and learning rate decreases ‘too fast’, consequently implying that cooperative strategies have not been explored enough during training. This is opposite to what the authors find with more sophisticated algorithms where collusive behaviours do not emerge. The conclusion is that the supra-competitive outcomes should be of concern for market authorities if and only if the algorithms employed are not sophisticated. Also Ref. [undefae] investigates when supra-competitive equilibria result necessarily from collusive strategies based on reward and punishment mechanisms to sustain cooperation. Unlike previous studies, their analysis critically suggests that both high prices and these strategic patterns actually emerge from a combination of simultaneous exploration and learning inertia intrinsic to reinforcement learning techniques, without any direct causal connection. They show that such forms of apparent collusion can arise even in memoryless settings and with myopic [^4] Agents with a low discount factor on future rewards, i.e. that value more immediate rewards over future possible ones. agents. Based on these findings, they caution against making straightforward analogies with human collusion and suggest addressing elevated algorithmic prices through tailored regulatory measures.
@@
There is no unanimous consensus on whether algorithmic collusion does –in fact– naturally emerge from the interaction of autonomous agents, and on the drivers that lead to supra-competitive equilibria. Especially, because of the definition of collusion555“Collusion is when firms use strategies that embody a reward-punishment scheme which rewards a firm for abiding by the supra-competitive outcome and punishes it for departing from it” as defined in [undefab]. that, according to [undef] is more suitable to a human to human cooperative behaviour rather than to an algorithmic driven phenomenon. Thus, a part of the literature points out at these phenomena as tacitly collusive equilibria while another stream interprets these results as not necessarily collusive even if the strategies employed by the agents might suggest so. This implies that applying existing regulatory solutions to markets where prices are set through reinforcement learning algorithms may miss the point, and thus interventions by the regulator on such markets might eventually be inadequate. In this fragmented landscape, literature on collusion in financial markets, to the best of our knowledge, has just started to develop. In fact, for what concerns financial markets, the problem becomes more involved since the actors in the market do not directly set the price in a ‘one sided’ way as is the case for production economies that are studied in economic literature. Among the many contributions that study the emergence of collusive behaviours in financial markets, in [undefar] it is shown how tacit collusion arises between market makers, modelled using deep RL, in a competitive market. In [undefn] the authors show how market making algorithms tacitly collude to extract rents, and this behaviour strictly depends on tick size and coarseness of price grid. Finally, in [undefs] the authors use a multi-agent deep RL algorithm to model market makers competing in the same market, the authors show how competing market makers learn how to adjust their quotes, giving rise to tacit collusion by setting spread levels strictly above the competitive equilibrium level.
To the best of our knowledge no prior work did consider supra-competitive, possibly tacit collusive, equilibria in optimal execution problems. As discussed in the Conclusions section, the existence of such equilibria might indeed imply a loss of welfare for other market participants. Therefore, studying the behaviour of learning agents opens for rather new perspectives on how collusion eventually arises through learning. Finally, studying supra-competitive equilibria and potential collusion for this class of problems is thus instrumental to understanding, and eventually, easing the regulatory tasks of market authorities when it comes to assess possible informational distortions of assets prices which, in turn, might affect the correct functioning of financial markets, resulting in welfare loss for market participants. This paper aims at tackling these points, while enlarging the analysis on these equilibria to the optimal execution problem.
@@
The paper is organised as follows: Section 2 introduces the theoretical setting of the market impact game, Section 3 introduces the DDQL algorithm for the multi-agent optimal execution problem, Section 4 discusses the results for different parameter settings, in Section 5 we critically discuss whether the observed results are tacit collusion or supra-competitive outcomes and finally Section 6 provides conclusions and outlines further research directions.
@@ 2 Market impact game setting
The setup of our framework is based on the seminal Almgren-Chriss model first introduced in [undefb] for optimal execution. In this setting, a single agent wants to liquidate an initial inventory of $q_{0}$ shares within a time window $[0,T]$, which, in the discrete-time setting, is divided into $N$ equal time increments of length $\tau=T/N$. The main assumption of the model is that the mid-price (or the efficient price) evolves according to a random walk with a drift depending on the traded quantity. Moreover, the price obtained by the agent differs from the mid-price by a quantity that depends on the volume traded in the interval. More formally, let $S_{t}$ and $\tilde{S}_{t}$ be the mid-price and the price received by the agent at time $t$, and let $v_{t}$ be the number of shares traded by the agent in the same interval; then the dynamics is given by
@@
$\begin{split}{S}_{t}&=S_{t-1}-{p}\left({v_{t}}/{\tau}\right)\tau+\sigma\tau^{\frac{1}{2}}\xi_{t}\\ \tilde{S_{t}}&={S}_{t-1}-h\left({v_{t}}/{\tau}\right){\quad.}\end{split}$ (1)
@@
$S_{t}$ evolves because of a diffusion part $\xi_{t}\sim\mathcal{N}(0,1)$ multiplied by the price volatility $\sigma$ and a drift part, termed permanent impact, ${p}\left({v_{t}}/{\tau}\right)$, assumed to be linear and constant: ${p}\left({v_{t}}/{\tau}\right)=\kappa{v_{t}}/{\tau}$. The price $\tilde{S}_{t}$ received by the agent is equal to the mid-price $S_{t}$ but impacted by a temporary impact term also assumed to be linear and constant in time: $h(v_{t}/{\tau})=\alpha{v_{t}}/{\tau}$.
These two impacts are the classic components of market impact. In fact, they represent the adverse movement of the asset price generated by trades. Permanent impact reflects permanent changes in the asset’s fundamental price process, while the temporary impact should reflect some temporary market imbalances. The former is cumulatively assumed to impact the asset price at least until the end of the execution window, while the latter vanishes as soon as the trade has been performed.
@@
The aim of the agent is to unwind their initial portfolio maximising the cash generated by their trading over the $N$ time steps. This objective can be rewritten in terms of Implementation Shortfall (IS) [^6] Notice that the IS could, in principle, take negative values as well. But in the absence of exogenous price drift and linear impact functions, as in our case, its expected value is positive. that is defined in the single agent case as:
$IS(\vec{v})=S_{0}q_{0}-\sum^{N}_{t=1}\tilde{S}_{t}v_{t}{\quad,}$ (2)
where $\vec{v}=(v_{1},...,v_{N})$ is the vector that contains the positive quantities sold at each time step, i.e. $v_{t}\geq 0,\penalty 10000\ \forall\penalty 10000\ t\in[1,N]$ and $S_{0}$ is the mid-price at the beginning of the trading window. The optimisation problem faced by the agent can be written as
$\min_{\vec{v}}\,\,\mathbb{E}[IS(\vec{v})]{+}{\lambda}\mathbb{V}[IS(\vec{v})]\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ s.t.\penalty 10000\ \penalty 10000\ \sum_{t=1}^{N}v_{t}=q_{0}{\quad,}$ (3)
@@
where $\lambda$ is the risk aversion parameter of the agent, $\mathbb{E}[\cdot]$ and $\mathbb{V}[\cdot]$ refer to the expected value and the variance of the IS functional respectively. Under the linearity assumption of the two impacts, the problem can be easily solved analytically. In the following, we are going to consider risk neutral agents ($\lambda=0$) and in this case the optimal trading schedule is the TWAP, i.e. $v_{t}=q_{0}/N$, where the trading velocity is constant.
@@
Now we consider two agents selling an initial inventory $q_{0}$ shares within the same time window time window $[0,T]$. The traded quantities by the two agents in the time step $t$ are indicated with $v^{(1)}_{t}$ and $v^{(2)}_{t}$ and $V_{t}=v^{(1)}_{t}+v^{(2)}_{t}$ is the total quantity traded in the same interval. The equations for the dynamics in discrete time are
@@
$\begin{split}{S}_{t}&=S_{t-1}-{p}\left({V_{t}}/{\tau}\right)\tau+\sigma\tau^{\frac{1}{2}}\xi_{t}\\ \penalty 10000\ \penalty 10000\ \tilde{S_{t}}^{(k)}&={S}_{t-1}-h\left({v^{(k)}_{t}}/{\tau}\right)\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ k=1,2{\quad;}\end{split}$ (4)
@@
i.e. the mid-price is affected by the total traded volume $V_{t}$, while the price received by each agents depends on the quantity they trade. Notice that, according to [undefam] the temporary impact depends only on the volume traded by the agent singularly and not on the volume traded by the other agent. This corresponds to the case where the asset is relatively liquid. Although it would be interesting to extend the model to the case of a temporary impact that depends on both agents’ traded volumes, in the present work we will consider the case in Eq. (4). Since more than one agent are optimising their trading volumes and the cash received depends on the trading activity of the other agent, the natural setting to solve this problem is the one of game theory. Since each agent is not directly aware of the selling activity of the other agent and the two agents interact through their impact on the price process, the resulting problem is an open-loop game. The existence and uniqueness of a Nash equilibrium in such a symmetric open-loop game has been studied in [undefam] in the case of n different players and linear impact functions. First of all, they defined the Nash equilibrium, in continuous time and considering the vector of levels of inventory held at each time-step $t$, i.e. $\vec{q}^{\penalty 10000\ (k)}$, as:
Definition 1 (Nash equilibrium [undefam]).
@@
In an $n$ players game, where $n\in\mathbb{N}$, with $q^{(1)}_{0},\dots,q^{(n)}_{0}\in\mathbb{R}$ initial inventory holdings and $\lambda_{1},\dots,\lambda_{n}$ non-negative coefficients of risk aversion. A Nash equilibrium for mean-variance optimisation as in Eq. (3) is a collection of $\mathbf{q^{*}}=\{\vec{q}^{\,(1)^{*}},\dots,\vec{q}^{\,(n)^{*}}\}$ inventory holdings such that for each $k\in n,\penalty 10000\ {\vec{q}^{\penalty 10000\ (k)}}\in\mathcal{A}_{\text{det}}$ the mean variance functional is minimised:
$\mathbb{E}[IS(\vec{q}^{\,(k)}|\mathbf{q^{\,(-k)^{*}}})]+{\lambda_{k}}\mathbb{V}[IS(\vec{q}^{\,(k)}|\mathbf{{q}^{\,(-k)^{*}}})]$
@@
for each agent kk, [^7] Where $\mathcal{A}_{\text{det}}$ is the set of all admissible deterministic trading strategies, for a more formal definition see [undefam]. where $\mathbf{{q}^{(-k)^{*}}}$ are the strategies of the other players minus the strategy of the $k$-agent that is being considered.
The set of deterministic strategies, $\mathcal{A}_{\text{det}}$, in the definition is the set of the strategies that is not adapted to the mid-price level. Thus, it does not depend on the realisation of the noise $\xi_{t}$ in the dynamics of the price process.
@@
Ref. [undefam] proves that there exists a unique Nash equilibrium for the mean-variance optimisation problem. The unique Nash equilibrium strategy $q^{(k)^{*}}_{t}$, the Nash remaining inventory of agent $k$ at time $t$, for $n$ players is given as a solution to the second-order system of differential equations: